A particle with charge 12.5muC and mass 2.9*10^-5k
Answer #1

The magnetic force is given by:

F = qv\times B

Sin the movement is perpendicular to the magnetic field:

F = qvBsin(90)

This force is equal to the rotational movement:

F = m\frac{v^{2}}{r}

So:

qvBsin(90)= m\frac{v^{2}}{r}

v = \frac{qBr}{m}

v = \frac{(12.5x10^{-6}C)(1.02T)(22.1m)}{2.9x10^{-5}kg}

v = 9.72m/s

We also know that the angular speed is given by:

\omega = \frac{v}{r} = \frac{\theta }{t}

Where v is speed, r is radius, theta is angular displacement and t is time. And if it has to complete an orbit, the displacement is 2pi. So:

t = \frac{\theta r}{v}

t = \frac{2\pi (22.1m)}{9.72m/s}

t = 14.29s

ANSWER t = 14.3s

ANSWER is D.

I hope it helps!!